OpenStudy (anonymous):

If a truck has ten times the mass of a car and the two vehicles are pushed with an equal force, you would expect the acceleration of the truck to be A. 10 times that of the car. B. 100 times that of the car. C. 1/100 times that of the car. D. 1/10 times that of the car.

OpenStudy (anonymous):

I chose A...

OpenStudy (anonymous):

Well F=ma. rearrange for "a" which is a=f/m. if you increase the mass by X10 it will be 1/10 the speed of the car

OpenStudy (anonymous):

which is D

OpenStudy (anonymous):

It will be D. since F=ma, f(car)=ma F(truck)=10ma' F(car)=F(truck) ma=10ma' a/a'=10 a'=a/10 so the accelertion of truck is 1/10 of the car

OpenStudy (anonymous):

Thank you guys

OpenStudy (jtfan2000):

Agreed with A